An automobile accelerates from rest at 1 ft/s
1
Expert's answer
2021-07-19T03:15:50-0400

If we were to calculate the distance covered, the following will be the steps.

"First \space distance \\
v=u+at \\
v=0+30*1\\
v=30ft/s\\
Distance, s= ut+\frac{1}{2}at^2=0.5*900= 450 ft\\
\\
Second \space distance \\

Distance, s= v*t=30*120= 3600 ft\\

Third \space distance \\
v=u+at \\
0=30+a*15\\
a=2ft/s^2\\
Distance, v^2=u^2+2ast\\
0=900-2*2*s \implies s= 225 ft\\
Total = 450+3600+225=4275ft=1303.02 m"